4. The order calls for Nipride 50mg in 1 liter D5W to be infused at 0.3mcg/kg/minute. The patient weighs 209 pounds. What is the flow rate in mL/hour?
Answer #1

209 pounds = 95 kgs

0.3mcg \times 95 = 28.5 mcg/minute

28.5mcg \times 60 = 1710 mcg/hour = 1.71 mg/hr

The required medication = required dose \times available volume \div available dose

                                        = 1.71 mg \times 1000 ml \div 50 mg

                                        = 34.2 ml

The flow rate is 34.2 ml/hour